OpenStudy (ajanijones):

help i give medals Find the exact circumference. d = 19.7 cm

OpenStudy (ajanijones):

Plzz help me.... Lol im so sleepy i cant think

OpenStudy (disco619):

Circumference Formula: \[2 \pi r\]

OpenStudy (disco619):

2 multiplied by Pi multiplied by Radius

OpenStudy (ajanijones):

Ok.

OpenStudy (disco619):

If Diameter is 19.7 then radius would be it's exact half.

OpenStudy (ajanijones):

so the answer is 9.85

OpenStudy (disco619):

So \[19.7 \div 2\] \[9.85\] That is the Radius Now we will put this in the formula: \[2 \times \pi \times r\] \[2 \times \pi \times 9.85\] Now you can solve this easily im sure.

OpenStudy (ajanijones):

so it 9.85cm or 9.85πcm

OpenStudy (agent0smith):

d = 19.7 cm Multiply that number by pi. Circumference = pi times diameter That's it.

OpenStudy (agent0smith):

\[\large C = 2 \pi r\]or\[\large C = \pi D\]where D is diameter

OpenStudy (ajanijones):

Why do you follow me every where agent)smith

OpenStudy (disco619):

That makes more sense. You can directly multiply the diameter with Pi

OpenStudy (ajanijones):

Ok so my answer is 9.85cm

OpenStudy (ajanijones):

I did and i gota 9.85

OpenStudy (disco619):

Multiply the diameter with Pi

OpenStudy (agent0smith):

\[\large C = \pi \times D\]plug in the D which is 19.7 and you are done.

OpenStudy (ajanijones):

I got 19.70

OpenStudy (ajanijones):

so its 19.7 not 19.7π

OpenStudy (agent0smith):

\[\large C =? \pi\]replace the question mark and you're done.

OpenStudy (agent0smith):

Put 19.7 in the place of the ?

OpenStudy (ajanijones):

no my answers are A. 9.85 cm B. 19.7 cm C. 9.85π cm D. 19.7π cm

OpenStudy (disco619):

Ok listen... If you have been given the radius then u use this to find circumference: \[2 \times \pi \times r\] But you have the diameter 19.7 so simply do this to find circumference: \[d \times \pi\] Which means: \[19.7 \times \pi\]

OpenStudy (agent0smith):

We've basically given you the answer now...

OpenStudy (ajanijones):

I know its 19.7π

OpenStudy (disco619):

Yep

OpenStudy (ajanijones):

Yea thxx i just needed the explanation. thanks guys

OpenStudy (rebeccapink6398):

r u in k12? did u get it right???????????????/